OpenStudy (anonymous):

I need to find the inverse function of (square root of x) / ((square root of x) - 3).

OpenStudy (anonymous):

\[\sqrt{x}/\sqrt{x} -3\]

OpenStudy (anonymous):

The easiest way to do it is switch the \(x\)'s and \(y\)'s in \[y = \frac{\sqrt{x}}{\sqrt{x}-3}\] and then solve for y.

OpenStudy (anonymous):

I tried that, but I wasn't able to solve for y. >-<